Apply the technique of ANOVA to the following data showing the yields
of 3 varieties of a crop each from 4 blocks, and test whether the mean
yields of the varieties are equal or not. Also test the equality of the block
means.
Varieties Blocks

I II III IV
A 4 8 6 8
B 5 5 7 8
C 6 7 9 5
1
Expert's answer
2020-11-18T19:05:25-0500

The following table is obtained:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & A & B & C \\ \hline
 & 4 & 5 & 6 \\
 & 8 & 5 & 7 \\
& 6 & 7 & 9 \\
& 8 & 8 & 5 \\
 \hdashline
 Sum= & 26 & 25 & 27 \\
 \hdashline
 Average= & 6.5 & 6.25 & 6.75 \\
 \hdashline
 \sum_iX_{ij}^2= & 180 & 163 & 191 \\
 \hdashline
 St.Dev.= & 1.915 & 1.5 & 1.708 \\
 \hdashline
 SS= & 11 & 6.75 & 8.75 \\
 \hdashline
 n= & 4 & 4 & 4 \\
 \hdashline
\end{array}"


The total sample size is "N=12." Therefore, the total degrees of freedom are:


"df=12-1=11"

The between-groups degrees of freedom are "df_{between}=3-1=2," and the within-groups degrees of freedom are:


"df_{within}=df_{total}-df_{between}=11-2=9"

We need to compute the total sum of values and the grand mean. The following is obtained


"\sum_{\mathclap{ i, j}} X_{ij}=25+26+27=78"

Also, the sum of squared values is


"\sum_{\mathclap{ i, j}} X_{ij}^2=180+163+191=534"


The total sum of squares is computed as follows


"SS_{total}=\sum_{\mathclap{ i, j}} X_{ij}^2-\dfrac{1}{N}(\sum_{\mathclap{ i, j}} X_{ij})^2="

"=534-\dfrac{(78)^2}{12}=27"

The within sum of squares is computed as shown in the calculation below:


"SS_{within}=\sum SS_{withingroups}=11+6.75+8.75=26.5"

The between sum of squares is computed directly as shown in the calculation below:

The between sum of squares is computed directly as shown in the calculation below:

Now that sum of squares are computed, we can proceed with computing the mean sum of squares:


"MS_{between}=\dfrac{SS_{between}}{df_{between}}=\dfrac{0.5}{2}=0.25"

"MS_{within}=\dfrac{SS_{within}}{df_{within}}=\dfrac{26.5}{9}=2.944"

Finally, with having already calculated the mean sum of squares, the F-statistic is computed as follows:


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{0.25}{2.944}=0.085"

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2=\mu_3"

"H_1:" Not all means are equal

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.


Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df_1=2" and "df_2=9," therefore, the rejection region for this F-test is "R=\{F:F>F_c=4.2565\}"

Test Statistics


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{0.25}{2.944}=0.085"

Since it is observed that "F=0.085<4.2565=F_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 3 population means are equal, at the "\alpha=0.05" significance level.

Using the P-value approach:

The p-value is "p=0.919241," and since "p=0.919241>0.05=\alpha," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 3 population means are equal, at the "\alpha=0.05" significance level.



The following table is obtained:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c: c}
 & I & II & III & IV \\ \hline
 & 4 & 8 & 6 & 8 \\
 & 5 & 5 & 7 & 8 \\
& 6 & 7 & 9 & 5 \\
 \hdashline
 Sum= & 15 & 20 & 22 & 21 \\
 \hdashline
 Average= & 5 & 6.667 & 7.333 & 7 \\
 \hdashline
 \sum_iX_{ij}^2= & 77 & 138 & 166 & 153 \\
 \hdashline
 St.Dev.= & 1 & 1.528 & 1.528 & 1.732 \\
 \hdashline
 SS= & 2 & 4.667 & 4.667 & 6 \\
 \hdashline
 n= & 3 & 3 & 3 & 3 \\
 \hdashline
\end{array}"


The total sample size is "N=12." Therefore, the total degrees of freedom are:


"df=12-1=11"

The between-groups degrees of freedom are "df_{between}=4-1=3," and the within-groups degrees of freedom are:


"df_{within}=df_{total}-df_{between}=11-3=8"

We need to compute the total sum of values and the grand mean. The following is obtained


"\sum_{\mathclap{ i, j}} X_{ij}=15+20+22+21=78"

Also, the sum of squared values is


"\sum_{\mathclap{ i, j}} X_{ij}^2=77+138+166+153=534"


The total sum of squares is computed as follows


"SS_{total}=\sum_{\mathclap{ i, j}} X_{ij}^2-\dfrac{1}{N}(\sum_{\mathclap{ i, j}} X_{ij})^2="

"=534-\dfrac{(78)^2}{12}=27"

The within sum of squares is computed as shown in the calculation below:


"SS_{within}=\sum SS_{withingroups}=""=2+4.667+4.667+6=17.333"

The between sum of squares is computed directly as shown in the calculation below:

The between sum of squares is computed directly as shown in the calculation below:

Now that sum of squares are computed, we can proceed with computing the mean sum of squares:


"MS_{between}=\dfrac{SS_{between}}{df_{between}}=\dfrac{9.667}{3}=3.222"

"MS_{within}=\dfrac{SS_{within}}{df_{within}}=\dfrac{17.333}{8}=2.167"

Finally, with having already calculated the mean sum of squares, the F-statistic is computed as follows:


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{3.222}{2.167}=1.487"

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2=\mu_3=\mu_4"

"H_1:" Not all means are equal

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.


Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df_1=3" and "df_2=8," therefore, the rejection region for this F-test is "R=\{F:F>F_c=4.0662\}"

Test Statistics


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{3.222}{2.167}=1.487"

Since it is observed that "F=1.487<4.0662=F_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 4 population means are equal, at the "\alpha=0.05" significance level.

Using the P-value approach:

The p-value is "p=0.290001," and since "p=0.290001>0.05=\alpha," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 4 population means are equal, at the "\alpha=0.05" significance level.